OpenStudy (anonymous):

What is the pressure of a system containing acetone at the temperature 34.3°C if the starting pressure and temperature are 1.03 atm and 23.8°C, respectively? given: Acetone normal boiling temp: 56.1 C Heat of vaporization: 30.3 (kj/mol)

OpenStudy (anonymous):

I know to solve I need Clausius-Clapeyron equation but I don't know how to use it.

OpenStudy (wolfe8):

Wouldn't you use P1/T1=P2/T2?

OpenStudy (aaronq):

P1/T1=P2/T2 is used when the substance is gaseous, at that temp, acetone is still a liquid, but you're finding the pressure caused by the vapour. clasius-clayperyon ln(P1/P2)=H(vap)/R*(1/T2-1/T1) plug in all the numbers: ln(1.03 atm/P2)=(30300 J/mol)/8.314 J/kmol)*(1/(273.15+34.3)-1/(273.15+23.8))

OpenStudy (wolfe8):

Ah my bad. I didn't realize this involves phase change.

OpenStudy (aaronq):

no worries, you live and learn